a 10% acid solution is to be mixed with a 50% solution to get 200 litres of a 20% solution. how many litres of each solution should be used?
Answer by NancyLam(35)

a 10% acid solution is to be mixed with a 50% solution to get 200 litres of a 20% solution. how many litres of each solution should be used?
let x = 10%
let y = 50%
x + y = 200
y= 200 -x
0.10x+0.50y = 200* 0.20
substitution : y= 200 -x
0.10x+0.50(200-x) = 200* 0.20
0.10x + 100- 0.50x = 40
-0.4x = 40 -100
x= -60/-0.4
x=15
substitution : x = 15 find y
15 + y = 200
y = 200-15
y = 185
10% acid solution is should be used 15 litres
50% acid solution is should be used 185 litres